ax^2-2x+9=0 Find all the possible values of a and find the equation has only one root.

ax^2-2x+9=0
Find all the possible values of a and find the equation has only one root.
 

Grade:12th pass

2 Answers

Arun
4 years ago
Dear student
 
if equation has only one root then its discriminant is 0
 
4 – 36 a = 0
 
a = 1/9
 
Hope it helps
Aditya Gupta
4 years ago
hello student. note that aruns answer is incomplete.
correct answer is:
if ax^2-2x+9=0 is a quad eqn, then discriminant is 0
4 – 36 a = 0
a = 1/9.
but if it is not a quadratic, then a= 0 in which case it becomes – 2x+9= 0 or x= 9/2, which is only a single root.
so, possible values of a are 1/9 and 0.
kindly approve :)